# Quantum Blackwell’s Ordering and Differential Privacy

### v1

@@ II Notations, Definitions and Facts
We use $\mathcal{H}$ to denote a finite-dimensional Hilbert space and we denote its dimension with $\left|\mathcal{H}\right|$, $\mathcal{D}(\mathcal{H})$ to represent the set of all state density matrices acting on $\mathcal{H}$. For any quantum state $\rho\in\mathcal{D}(\mathcal{H})$, we define $\mbox{supp}(\rho):=\mbox{Span}\{|i\rangle:\lambda_{i}>0\},$ where $\left\{\lambda_{i}\right\}$ represent non-zero eigenvalues of $\rho$. For any finite and non-empty set $\mathcal{X}$, we denote $\mathcal{P}(\mathcal{X})$ as the set of all probability distributions over $\mathcal{X}$. Similarly, for any distribution $P\in\mathcal{P}(\mathcal{X})$, $\mbox{supp}(P):=\{x\in\mathcal{X}:P(x)>0\}.$ For any sequence $x^{n}:(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}$, a permutation $\pi:\mathcal{X}^{n}\to\mathcal{X}^{n},$ is a map such that $\pi\left(x^{n}\right):=(x_{\pi^{-1}(1)},\cdots,x_{\pi^{-1}(n)}).$ Let $S_{n}$ be the permutation group containing all permutations of length $n$ and it is also known as $n$-th symmetric group.
Definition 1.
(Type and Set of all types) A vector of integers ${\bf f}:=\left(f_{1},\cdots,f_{d}\right)$ (where $d>0$) is called a type of size $n$ and length $d$ if $\forall i\in[d],f_{i}\geq 0$ and $\sum_{i}f_{i}=n.$ $T^{n}_{d}$ is denoted as the set of all distinct types of size $n$ and length $d$.
Definition 2 (Frequency type classes).
For a non-empty finite set $\mathcal{X}$ and an integer $n>0$, given a type function ${\bf f}:=(f_{1},\cdots,f_{\left|\mathcal{X}\right|}])$ ( such that $\sum_{i\in\mathcal{X}}f_{i}=n$), we define a set $T_{{\bf f}}\subset\mathcal{X}^{n}$ to be the frequency type class or type set corresponding to $\bf f$} as follows,
$T_{\bf f}:=\left\{(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}:\frac{\left\lvert\left\{k:x_{k}=i\right\}\right\rvert}{n}=\bar{f}(i),\forall i\in\mathcal{X}\right\},$
where $\bar{f}(i):=\frac{f_{i}}{n}$ for all $i\in\mathcal{X}$.
Definition 3 ([6]).
A pair of $n$-length (where $n\in\mathbb{N}$ ) sequences $x^{n}:=(x_{1},\cdots,x_{n}),\tilde{x}^{n}:=(\tilde{x}_{1},\cdots,\tilde{x}_{n})\in\mathcal{X}^{n}$ (where $\left|\mathcal{X}\right|>0$) is said to be $k$-neighbors, represented by $x^{n}\overset{k}{\sim}\tilde{x}^{n}$ if,
$k=\frac{1}{2}\sum_{a\in\mathcal{X}}\left|f(a|x^{n})-f(a|\tilde{x}^{n})\right|,$ (1)
where, $f(a|x^{n})$ denotes the number of appearance of the alphabet $a\in\mathcal{X}$ in the sequence $x^{n}$.
Further, for any sequence $x^{n}$ and its some permutation $\tilde{x}^{n}$, we denote $x^{n}\overset{0}{\sim}\tilde{x}^{n}$.
Definition 4 (Minimun type-II error).
@@
$\beta(\alpha,P_{1},P_{2}):=\min_{\begin{subarray}{c}\phi:\mathcal{X}\to[0,1]\\ \sum_{x\in\mathcal{X}}P_{1}(x)\phi(x)\leq\alpha\end{subarray}}1-\sum_{x\in\mathcal{X}}P_{2}(x)\phi(x).$ (2)
@@
Definition 7 (Quantum Instrument [22]).
A quantum instrument consists of a collection $\{\mathcal{E}_{j}\}$ of completely positive, trace non-increasing maps such that the sum map $\sum_{j}\mathcal{E}_{j}$ is trace preserving. Let $\{|j\rangle\}_{j}$ be an orthonormal basis for a Hilbert space $\mathcal{H}_{J}$. The action of a quantum instrument on a density operator $\rho\in\mathcal{D}(\mathcal{H})$ is the following quantum channel, which features a quantum and classical output:
$\rho\mapsto\sum_{j}\mathcal{E}_{j}(\rho)\otimes|j\rangle\langle j|_{J}.$ (4)
Definition 8 (Smooth Max-Relative Entropy).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and $\varepsilon\in[0,1)$, the smooth max-relative entropy of $\rho$ with respect to $\sigma$ is defined as
$D_{\max}^{\varepsilon}(\rho\|\sigma):=\min_{\tilde{\rho}\in\mathcal{B}^{\varepsilon}(\rho)}D_{\max}(\tilde{\rho}\|\sigma),$ (5)
where $D_{\max}(\tilde{\rho}\|\sigma):=\inf\{\lambda\in\mathbb{R}:\tilde{\rho}\leq 2^{\lambda}\sigma\}$ and $\mathcal{B}^{\varepsilon}(\rho):=\left\{\tilde{\rho}\in\mathcal{D}(\mathcal{H}_{A}):\frac{1}{2}\|\tilde{\rho}-\rho\|_{1}\leq\varepsilon\right\}$.
@@
$\mathbb{D}(P_{1}\|P_{2})\geq\mathbb{D}(P^{\mathcal{K}}_{1}\|P^{\mathcal{K}}_{2}),\forall P_{1},P_{2}\in\mathcal{P}(\mathcal{X}),$ (6)
where for each $i=1,2$, $P^{\mathcal{K}}_{i}(y):=\mathbb{E}_{X\sim P_{i}}\left[\mathcal{K}(y\mid X)\right]$.
@@
Definition 11.
Definition 12 (Petz Quantum Rényi Divergence [23]).
Consider $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and $\alpha\in[0,1]\cup(1,+\infty)$. Then, Petz Quantum Rényi Divergence of order $\alpha$ between $\rho$ and $\sigma$ is defined as follows,
$D_{\alpha}(\rho\|\sigma):=\begin{cases}\frac{1}{\alpha-1}\log\mathrm{Tr}\left[\rho^{\alpha}\sigma^{1-\alpha}\right],&{\mbox{if }}(\alpha<1\cap\rho\not\perp\sigma)\cup(\rho\ll\sigma),\\ +\infty,&\mbox{else}.\end{cases}$
Definition 13 (Quantum hockey stick divergence [24]).
@@
For any type ${\bf f}=(f_{1},\cdots,f_{d})$, the set $T_{\bf f}$ (see Definition 2) is a permutation invariant set, i.e., for every sequence $x^{n}:(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}$, under any permutation $\pi\in S_{n},$ $\pi\left(x^{n}\right)\in T_{\bf f}$ where $\pi_{n}\left(x^{n}\right):=(x_{\pi^{-1}(1)},\cdots,x_{\pi^{-1}(n)}).$
Fact 2.
[25] $T^{n}_{d}$ (See Definition 1) satisfies the following upper-bound :
$\left|\Lambda^{n}_{d}\right|\leq\left|T^{n}_{d}\right|\leq(n+1)^{d-1}.$ (9)
Fact 3.
@@
$E_{\gamma}(P^{\prime}\|Q^{\prime})$ $\leq E_{\gamma}(P\|Q)+{\left\lVert P^{\prime}-P\right\rVert}_{1}+\gamma{\left\lVert Q^{\prime}-Q\right\rVert}_{1}.$ (10)
@@
$E_{\gamma}(P^{\prime}\|Q^{\prime})$ $=\max_{\begin{subarray}{c}0\leq\Lambda(x)\leq 1,\\ x\in\mathcal{X}\end{subarray}}\sum_{x\in\mathcal{X}}\Lambda(x)(P^{\prime}(x)-\gamma Q^{\prime}(x))$
$\leq\max_{\begin{subarray}{c}0\leq\Lambda(x)\leq 1,\\ x\in\mathcal{X}\end{subarray}}\sum_{x\in\mathcal{X}}\Lambda(x)(P(x)-\gamma Q(x))+\sum_{x\in\mathcal{X}}|P^{\prime}(x)-P(x)|+\gamma\sum_{x\in\mathcal{X}}|Q^{\prime}(x)-Q(x)|$
$\leq E_{\gamma}(P\|Q)+{\left\lVert P^{\prime}-P\right\rVert}_{1}+\gamma{\left\lVert Q^{\prime}-Q\right\rVert}_{1}.$
This proves Fact 3.
Fact 4 (Hölder’s inequality).
@@
Further, if $p=1$ and $q=\infty$, then the right-hand side of (11) can be interpreted as follows,
@@
$D(P\|Q)=\int_{0}^{\infty}\left(\frac{1}{\gamma}E_{\gamma}(P\|Q)+\frac{1}{\gamma^{2}}E_{\gamma}(Q\|P)\right)d\gamma,$ (13)
@@
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, the quantum relative entropy $D(\rho\|\sigma)$ has the following integral representation in terms of quantum hockey stick divergence,
@@
Fact 7.
@@
$\min_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]\leq\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma]=\min_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]=\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma].$ (15)
@@
$\mathrm{Tr}[(\mathbb{I}-\hat{\Lambda})\sigma]$ $=\delta\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]$
$\overset{a}{<}\beta^{\prime},$
where $a$ follows from the fact that $\delta<1$ and $\beta^{\prime}=\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]$. Thus, for any POVM $\{\tilde{\Lambda},\mathbb{I}-\tilde{\Lambda}\}$, such that $\mathrm{Tr}[\tilde{\Lambda}\rho]<\alpha$, there exists a POVM $\{\hat{\Lambda},\mathbb{I}-\hat{\Lambda}\}$ such that $\mathrm{Tr}[\tilde{\Lambda}\rho]=\alpha$ and $\mathrm{Tr}[(\mathbb{I}-\hat{\Lambda})\sigma]<\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]$. Therefore, the optimum $\Lambda$ in the LHS of (15) satisfies $\mathrm{Tr}[{\Lambda}\rho]=\alpha$. This proves Fact 7.
Fact 8 ([24]).
@@
$E_{\gamma}(\rho\|\sigma)=\mathrm{Tr}|\Lambda(\rho-\gamma\sigma)|_{+},$ (16)
@@
Fact 9.
@@
Proof of Fact 9 follows directly from the proof of Fact 7.
Fact 10 (Reverse quantum data processing inequality).
Consider $\rho_{1},\rho_{2}\in\mathcal{D}(\mathcal{H}_{A})$ and $\sigma_{1},\sigma_{2}\in\mathcal{D}(\mathcal{H}_{B})$. Then, if $\forall\gamma\geq 0$, $E_{\gamma}(\rho_{1}\|\rho_{2})\geq E_{\gamma}(\sigma_{1},\sigma_{2})$, then there exists a completely positive map $\mathcal{T}:\mathcal{H}_{A}\to\mathcal{H}_{B}$ such that $\sigma_{i}=\mathcal{T}(\rho_{i})$, for each $i\in\{1,2\}$.
Proof.
The proof of Fact 10 follows directly from Theorem 55 and Theorem $6$ of [27].
Fact 11 (Data-processing inequality of quantum relative entropy).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H})$, $D(\rho\|\sigma)$ the following holds,
$D(\rho\|\sigma)\geq D(\mathcal{E}(\rho)\|\mathcal{E}(\sigma)),$ (18)
where $\mathcal{E}$ is any completely positive and trace-preserving (CP-TP) map.
Fact 12 (Data-processing inequality for quantum hypothesis testing divergence [28]).
Consider any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and any CP-TP map $\mathcal{N}:\mathcal{H}_{A}\to\mathcal{H}_{B}$. Then for any $\alpha\in[0,1]$ the following satisfies,
$\mathcal{D}^{\alpha}_{H}(\rho\|\sigma)\geq\mathcal{D}^{\alpha}_{H}(\left(\mathcal{N}(\rho)\|\mathcal{N}(\sigma)\right).$ (19)
Fact 13 (Data-processing inequality of Petz quantum Rényi divergence [23]).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H})$ and $\forall\alpha\in[0,1)\cup(1,2]$, $D_{\alpha}(\rho\|\sigma)$ satisfies the following,
$D_{\alpha}(\rho\|\sigma)\geq D_{\alpha}(\mathcal{E}(\rho)\|\mathcal{E}(\sigma)),$ (20)
where $\mathcal{E}$ is any completely positive and trace-preserving (CP-TP) map.
Fact 14 (Data-processing inequality of quantum Hockey stick divergence [24, Lemma $4$]).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H})$ and $\forall k\geq 0$, $E_{k}(\rho\|\sigma)$ satisfies the following,
$E_{k}(\rho\|\sigma)\geq E_{k}(\mathcal{E}(\rho)\|\mathcal{E}(\sigma)),$ (21)
where $\mathcal{E}$ is any completely positive and trace-preserving (CP-TP) map.
Fact 15 (Monotonicity of SLD Fisher Information [29]).
Let $\rho_{\theta}$ be a differentiable family of quantum states and let $\mathcal{E}$ be a completely positive trace-preserving (CP-TP) map. Then, the SLD Fisher information is monotonic under $\mathcal{E}$:
$J_{\theta}(\rho_{\theta})\geq J_{\theta}(\mathcal{E}(\rho_{\theta})).$ (22)
Fact 16 (Quantum Reversed Pinsker inequality[30, theorem 2]).
For quantum states $\rho$ and $\sigma$, the following inequality holds:
$D(\rho\,\|\,\sigma)\leq\frac{2}{\lambda_{\min}(\sigma)}E_{1}(\rho\,\|\,\sigma)^{2}.$ (23)
where $\lambda_{\min}(\sigma)$ is the smallest non-zero eigenvalue of $\sigma$.
Fact 17.
Consider $\rho,\rho^{\prime},\sigma\in\mathcal{D}(\mathcal{H}_{A})$. Then, the following inequality holds:
$D(\rho\|\sigma)\leq D(\rho\|\rho^{\prime})+D_{\max}(\rho^{\prime}\|\sigma).$ (24)
Proof.
By the definition of the max-relative entropy, we have
$\rho^{\prime}\leq e^{D_{\max}(\rho^{\prime}\|\sigma)}\,\sigma.$
Equivalently,
$\sigma\geq e^{-D_{\max}(\rho^{\prime}\|\sigma)}\,\rho^{\prime}.$
Since the logarithm is operator monotone, this implies
$\log\sigma\succeq\log\rho^{\prime}-D_{\max}(\rho^{\prime}\|\sigma)\,\mathbb{I}.$
Multiplying both sides by $\rho$ and taking the trace (which reverses the inequality), we obtain
$\mathrm{Tr}[\rho\log\sigma]\leq-\mathrm{Tr}[\rho\log\rho^{\prime}]+D_{\max}(\rho^{\prime}\|\sigma).$
Adding $\mathrm{Tr}[\rho\log\rho]$ to both sides gives
$\mathrm{Tr}[\rho(\log\rho-\log\sigma)]\leq\mathrm{Tr}[\rho(\log\rho-\log\rho^{\prime})]+D_{\max}(\rho^{\prime}\|\sigma),$
which can be written as
$D(\rho\|\sigma)\leq D(\rho\|\rho^{\prime})+D_{\max}(\rho^{\prime}\|\sigma).$
This completes the proof of Fact 17.
Fact 18.

### v3

@@ II Notations, Definitions and Facts
We use $\mathcal{H}$ to denote a finite-dimensional Hilbert space and we denote its dimension with $\left|\mathcal{H}\right|$, $\mathcal{D}(\mathcal{H})$ to represent the set of all density matrices acting on $\mathcal{H}$. For any quantum state $\rho\in\mathcal{D}(\mathcal{H})$, we define $\mbox{supp}(\rho):=\mbox{Span}\{|i\rangle\mid\lambda_{i}>0\},$ where $\left\{\lambda_{i}\right\}$ represent non-zero eigenvalues of $\rho$. For any finite and non-empty set $\mathcal{X}$, we denote $\mathcal{P}(\mathcal{X})$ as the set of all probability distributions over $\mathcal{X}$. Similarly, for any distribution $P\in\mathcal{P}(\mathcal{X})$, we denote $\mbox{supp}(P):=\{x\in\mathcal{X}\mid P(x)>0\}.$
Definition 2 (Classical minimum type-II error).
@@
$\beta(\alpha|P_{1}\|P_{2}):=\min_{\begin{subarray}{c}\phi:\mathcal{X}\to[0,1]\\ \sum_{x\in\mathcal{X}}P_{1}(x)\phi(x)\leq\alpha\end{subarray}}1-\sum_{x\in\mathcal{X}}P_{2}(x)\phi(x).$ (2)
@@
Definition 5 (Classical alpha divergence).
For any pair of probability distributions $P,Q$ over a finite set $\mathcal{X}$, and for any $\alpha\in[0,1]$, the classical $\alpha$ divergence of order $\alpha$ is defined as
$D_{\alpha}(P\|Q):=\frac{1}{\alpha-1}\log\sum_{x\in\mathcal{X}}P(x)^{\alpha}Q(x)^{1-\alpha}.$ (4)
@@
$\mathbb{D}(P_{1}\|P_{2})\geq\mathbb{D}(\mathcal{K}(P_{1})\|\mathcal{K}(P_{2})),\forall P_{1},P_{2}\in\mathcal{P}(\mathcal{X}),$ (5)
where for each $i=1,2$, $\mathcal{K}(P_{i})(y):=\mathbb{E}_{X\sim P_{i}}\left[\mathcal{K}(y\mid X)\right]$.
Definition 7 (Quantum minimum type-II error).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, the minimum type-II error of a fixed order $\alpha\in[0,1]$ is defined as follows,
$\beta(\alpha|\rho\|\sigma):=\min_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]\leq\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma].$ (6)
Definition 8 (Quantum hypothesis testing divergence [32]).
Definition 9 (Quantum hockey stick divergence [30]).
Definition 10 ([19]).
Consider $\mathcal{F}$ be the set of functions $f:(0,\infty):\to\mathbb{R}$ that are convex and twice differentiable with $f(1)=0$. Then, for any $f\in\mathcal{F}$ and for any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we define the quantum $f$-divergence $D_{f}(\rho\|\sigma)$ as follows,
$D_{f}(\rho\|\sigma):=\int_{1}^{\infty}\left(f^{\prime\prime}(\gamma)E_{\gamma}(\rho\|\sigma)+\gamma^{-3}f^{\prime\prime}(\gamma^{-1})E_{\gamma}(\sigma\|\rho)\right)d\gamma.$ (9)
whenever, the integral is finite and $D_{f}(\rho\|\sigma):=+\infty$ otherwise.
Definition 11.
Definition 12 (Hockey Stick Rényi Divergence [19]).
Consider $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and $\alpha\in[0,1)\cup(1,+\infty)$. Then, hockey stick Rényi Divergence of order $\alpha$ between $\rho$ and $\sigma$ is defined as follows,
$D_{\alpha}(\rho\|\sigma)$ $:=\frac{1}{\alpha-1}\log\big(1+(\alpha-1)H_{\alpha}(\rho\|\sigma)\big),$ (10)
$H_{\alpha}(\rho\|\sigma)$ $:=\alpha\int_{1}^{\infty}\gamma^{\alpha-2}E_{\gamma}(\rho\|\sigma)+\gamma^{-\alpha-1}E_{\gamma}(\sigma\|\rho)\,\mathrm{d}\gamma,$ (11)
when a parameter $\alpha\in[0,1)\cup(1,+\infty)$ satisfies the condition $\alpha<1\cap\rho\not\perp\sigma$ or the condition $\rho\ll\sigma$.
Definition 13 (Measured Rényi Divergence [16, 19]).
Consider $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and $\alpha\in[0,1)\cup(1,+\infty)$. Then, measured Rényi Divergence of order $\alpha$ between $\rho$ and $\sigma$ is defined as follows,
$\check{D}_{\alpha}(\rho\|\sigma):=\sup_{M:{\rm POVM}}D_{\alpha}(P^{M}_{\rho}\|P^{M}_{\sigma}),$
where $P^{M}_{\rho}$ is the distribution given as $P^{M}_{\rho}(i):=\mathrm{Tr}\rho M_{i}$ for a POVM $M=\{M_{i}\}_{i}$.
@@
$E_{\gamma}(P^{\prime}\|Q^{\prime})$ $\leq E_{\gamma}(P\|Q)+\frac{1}{2}{\left\lVert P^{\prime}-P\right\rVert}_{1}+\frac{\gamma}{2}{\left\lVert Q^{\prime}-Q\right\rVert}_{1}.$ (13)
@@
$E_{\gamma}(P^{\prime}\|Q^{\prime})$ $\overset{(a)}{=}P^{\prime}(A)-\gamma Q^{\prime}(A)$
$\leq P(A)-\gamma Q(A)+P^{\prime}(A)-P(A)+\gamma(Q^{\prime}(A)-Q(A))$
$\leq\max_{B\subseteq\mathcal{X}}(P(B)-\gamma Q(B))+\max_{C\subseteq\mathcal{X}}(P^{\prime}(C)-P(C))+\gamma\max_{D\subseteq\mathcal{X}}(Q^{\prime}(D)-Q(D))$
$=E_{\gamma}(P\|Q)+\frac{1}{2}{\left\lVert P^{\prime}-P\right\rVert}_{1}+\frac{\gamma}{2}{\left\lVert Q^{\prime}-Q\right\rVert}_{1}.$
where in $(a)$, let $A$ be the maximal set for $E_{\gamma}(P^{\prime}\|Q^{\prime})$ according to Definition 4. This proves Fact 1.
Fact 2 (Hölder’s inequality).
@@
Further, for the case when $p=1$ and $q=\infty$, the right-hand side of (14) can be interpreted as follows,
@@
$D(P\|Q)=\int_{1}^{\infty}\left(\frac{1}{\gamma}E_{\gamma}(P\|Q)+\frac{1}{\gamma^{2}}E_{\gamma}(Q\|P)\right)d\gamma,$ (16)
@@
Fact 4.
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, for any $\alpha\in[0,1]$, the quantum hypothesis testing divergence $D^{\alpha}_{H}{(\rho\|\sigma)}$ of order $\alpha$ satisfies the following,
$D^{\alpha}_{H}{(\rho\|\sigma)}=-\log\beta(\alpha|\rho\|\sigma),$ (17)
where for any $\alpha\in[0,1]$, $\beta(\alpha|\rho\|\sigma)$ is defined in Definition 7.
Fact 5 ([19]).
For any $f\in\mathcal{F}$ (where $\mathcal{F}$ is defined in Definition 10) and for any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we have,
$D_{f}(\rho\|\sigma):=\int_{0}^{\infty}f^{\prime\prime}(\gamma)E_{\gamma}(\rho\|\sigma)d\gamma,$ (18)
where $D_{f}(\rho\|\sigma)$ is defined in Definition 10.
@@
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, and setting $f=x\log x$ in Definition 10, we have the following integral representation of the quantum relative entropy $D(\rho\|\sigma)$,
@@
Fact 7 ([19]).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we have the following relation between hockey-stick Rényi Divergence $D_{\alpha}(\rho\|\sigma)$ and measured Rényi Divergence $\check{D}_{\alpha}(\rho\|\sigma)$,
$\check{D}_{\alpha}(\rho\|\sigma)\leq D_{\alpha}(\rho\|\sigma),$ (20)
for a parameter $\alpha\in[0,1)\cup(1,+\infty)$.
Fact 8.
Let $\rho$ and $\sigma$ be two commuting quantum states on a Hilbert space $\mathcal{H}$ (i.e., $\rho,\sigma]=0$). For any order $\alpha\in(0,\infty)\setminus\{1\}$, the Hockey Stick Rényi divergence coincides with the Measured Rényi divergence:
$D_{\alpha}(\rho\|\sigma)=\check{D}_{\alpha}(\rho\|\sigma)=\frac{1}{\alpha-1}\log\left(\sum_{i}p_{i}^{\alpha}q_{i}^{1-\alpha}\right),$ (21)
where $\{p_{i}\}$ and $\{q_{i}\}$ are the eigenvalues of $\rho$ and $\sigma$ in their common eigen-basis.
Fact 9.
@@
$\beta(\alpha|\rho\|\sigma)=\min_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]\leq\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma]=\min_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]=\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma].$ (22)
@@
$\mathrm{Tr}[(\mathbb{I}-\hat{\Lambda})\sigma]=\delta\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]\overset{(a)}{<}\beta^{\prime},$
where $(a)$ follows from the fact that $\delta<1$ and $\beta^{\prime}=\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]$. Thus, for any POVM $\{\tilde{\Lambda},\mathbb{I}-\tilde{\Lambda}\}$, such that $\mathrm{Tr}[\tilde{\Lambda}\rho]<\alpha$, there exists a POVM $\{\hat{\Lambda},\mathbb{I}-\hat{\Lambda}\}$ such that $\mathrm{Tr}[\tilde{\Lambda}\rho]=\alpha$ and $\mathrm{Tr}[(\mathbb{I}-\hat{\Lambda})\sigma]<\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]$. Therefore, the optimum $\Lambda$ in the LHS of (22) satisfies $\mathrm{Tr}[{\Lambda}\rho]=\alpha$. This proves Fact 9.
Fact 10 ([30]).
@@
$E_{\gamma}(\rho\|\sigma)=\mathrm{Tr}|(\rho-\gamma\sigma)|_{+},$ (23)
@@
Fact 11 ([6, Lemma $3.1$]).
Let $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathbb{C}^{2})$ be commuting qubit quantum states and let $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathbb{C}^{d})$. Then the following two conditions are equivalent:
  1. 1.
    There exists a CP-TP map $\mathcal{E}:\mathcal{S}(\mathbb{C}^{2})\to\mathcal{S}(\mathbb{C}^{d})$ such that $\mathcal{E}(\rho_{1})=\rho_{2}$ and $\mathcal{E}(\sigma_{1})=\sigma_{2}$;
  2. 2.
    $D_{\max}(\rho_{1}\|\sigma_{1})\geq D_{\max}(\rho_{2}\|\sigma_{2})$ and $D_{\max}(\sigma_{1}\|\rho_{1})\geq D_{\max}(\sigma_{2}\|\rho_{2})$.
Fact 12.
@@
Proof of Fact 12 follows directly from the proof of Fact 9.
Fact 13.